# JaxRobotarium: Training and Deploying Multi-Robot Policies in 10 Minutes

### v2

@@ Abstract
Multi-agent reinforcement learning (MARL) has emerged as a promising solution for learning complex and scalable coordination behaviors in multi-robot systems. However, established MARL platforms (e.g., SMAC and MPE) lack robotics relevance and hardware deployment, leaving multi-robot learning researchers to develop bespoke environments and hardware testbeds dedicated to the development and evaluation of their individual contributions. The Multi-Agent RL Benchmark and Learning Environment for the Robotarium (MARBLER) is an exciting recent step in providing a standardized robotics-relevant platform for MARL, by bridging the Robotarium testbed with existing MARL software infrastructure. However, MARBLER lacks support for parallelization and GPU/TPU execution, making the platform prohibitively slow compared to modern MARL environments and hindering adoption. We contribute JaxRobotarium, a Jax-powered end-to-end simulation, learning, deployment, and benchmarking platform for the Robotarium. JaxRobotarium enables rapid training and deployment of multi-robot reinforcement learning (MRRL) policies with realistic robot dynamics and safety constraints, supporting both parallelization and hardware acceleration. Our generalizable learning interface provides an easy-to-use integration with SOTA MARL libraries (e.g., JaxMARL). In addition, JaxRobotarium includes eight standardized coordination scenarios, including four novel scenarios that bring established MARL benchmark tasks (e.g., RWARE and Level-Based Foraging) to a realistic robotics setting. We demonstrate that JaxRobotarium retains high simulation fidelity while achieving dramatic speedups over baseline (20x in training and 150x in simulation), and provides an open-access sim-to-real evaluation pipeline through the Robotarium testbed, accelerating and democratizing access to multi-robot learning research and evaluation. Our code is publicly available at https://github.com/GT-STAR-Lab/JaxRobotarium.
@@ 1 Introduction
Despite its potential, learning-based algorithms are seldom deployed and evaluated on physical multi-robot platforms. Even when physical deployments are undertaken, they tend to rely on bespoke and closed platforms and testbeds, limiting reproducibility and benchmarking. In stark contrast, the multi-agent reinforcement learning (MARL) research community outside of robotics benefits from standardized and structured benchmarks and simulation and learning environments (e.g., Multi-Agent Particle Environment (MPE) [9], StarCraft Multi-Agent Challenge (SMAC) [10], and JaxMARL [11]). However, these environments often abstract away robotics-relevant characteristics (e.g., robot dynamics and collision avoidance) and result in considerable sim-to-real gap [8], limiting their ability to faithfully evaluate learning approaches tailored to multi-robot systems.
MARBLER [12] is a recent framework that attempts to meet this critical need. MARBLER acts as a bridge between the standardized infrastructure (designed for MARL algorithms) and the Robotarium [13] (an open and remotely-accessible multi-robot testbed). However, MARBLER does not support parallelization and execution of MARL algorithms on GPU/TPU, resulting in prohibitively-slow training speeds that hinder sufficient training, adoption, and comprehensive evaluations.
@@
JaxRobotarium includes three key contributions. First, we contribute a new Jax-based simulator for the Robotarium that is designed from the ground-up and modeled after the existing Robotarium Python Simulator (RPS). Our new simulator supports parallelization and GPU/TPU execution, enabling dramatic improvements in efficiency and directly reducing the time and computational overhead typically associated with learning multi-robot coordination in realistic settings. Second, we contribute a new intuitive interface among a new simulator, the Robotarium, and existing MARL libraries, such as JaxMARL [11], allowing researchers to leverage and build upon existing algorithmic implementations with minimal setup. Third, we contribute a new unified benchmark for MRRL that currently supports 8 representative multi-robot coordination scenarios, with flexible tools to rapidly create new ones or customize the existing ones. The unique benefits of JaxRobotarium include:
@@
We validate the utility and benefits of JaxRobotarium through extensive efficiency evaluations, algorithm benchmarking, and comprehensive sim-to-real transfer experiments. We show that, compared with MARBLER [12], JaxRobotarium is able to simulate robot trajectories up to 150 times faster, and accelerates the training process of multi-robot policies up to 20 times faster. We benchmarked four MARL algorithms on the 8 scenarios provided by JaxRobotarium, revealing new insights into the algorithms’ utility across different multi-robot coordination scenarios. We systematically studied the sim-to-real gap of JaxRobotarium using data from over 200 real-world deployments of multi-robot policies that were entirely trained in simulation. We note that JaxRobotarium is not a panacea for sim-to-real gap. Rather, JaxRobotarium helps reduce sim-to-real gap by enabling rapid training and deployment. Our analysis reveals that while many task-algorithm combinations result in negligible performance differences when deployed on hardware, certain combinations of MARL algorithms and task characteristics exhibit noticeable performance differences. However, we show that such differences can be largely mitigated through simple randomization techniques such as noisy actions. We believe JaxRobotarium will serve as a powerful tool to advance the development, benchmarking, and sim-to-real transfer of learning algorithms aimed at the multi-robot systems community.
@@ 2 Related Work
The platforms to facilitate learning multi-robot coordination strategies should have the following characteristics: (i) support efficient training and evaluation of policies [11, 14]; (ii) have realistic simulations of real robots [8, 13], taking into account their dynamic constraints and the need for lower-level collision avoidance; (iii) have convenient interface with common MARL algorithms and benchmarking tasks [11, 12, 15]. Centering around these key requirements, we evaluate and compare available MRRL and MARL frameworks in Table 1. The Multi-Agent Particle Environment (MPE) [9] is a widely used framework for evaluating MARL algorithms on a diverse suite of 2D tasks. While MPE was enhanced by successors including VMAS [14], BenchMARL [15], and JaxMARL [11], none of them realistically reflects the dynamics of real robots, impeding straightforward sim2real evaluation of MRRL policies. To better evaluate MRRL policies, researchers must build task-specific test platforms [16, 17, 18] or reproduce previously proposed test beds, such as Cambridge RoboMaster [19]. As a contrastive line of efforts, the Robotarium [13] is a publicly available, multi-robot testbed that supports testing the coordination of up to 20 GRITSBots [20] robots. The Robotarium provides a simulator with implemented collision avoidance using control barrier functions [21] and offers sim-to-real evaluation through its hardware test bed. To bridge the gap between MARL and a real multi-robot testbed, [12] contributes MARBLER, which is an open platform that promotes the comprehensive evaluation of MRRL algorithms on the Robotarium. Unfortunately, MARBLER only runs on the CPU without parallelization, leading to prolonged training and tedious development cycles. JaxRobotarium builds upon the philosophy of MARBLER, marrying it with the computational efficiency offered by Jax and GPU parallelization.
@@
| MARL Platforms | ✓ | ✗ | ✗ | ✗ | ✓ | N/A |
| (MPE [9], SMAC [10], VMAS [14], BenchMARL [15], | ||||||
| JaxMARL [11], Google Research Football [22]) | ||||||
@@ 3 JaxRobotarium Platform
JaxRobotarium is designed to achieve the training speeds of Multi-Agent RL in Multi-Robot RL while maintaining the standardized, public, and free sim-to-real evaluation through the Robotarium. To this end, the JaxRobotarium platform consists of the following key components,
@@
A Jax redesign of the Robotarium Python Simulator (RPS), compatible with parallelization and GPU/TPU execution.
@@
A user-friendly deployment pipeline for sim-to-real evaluation on the Robotarium.
These components culminate in up to a 20x speedup over MARBLER. We dedicate the remainder of this paper to discussing and evaluating each component. A full overview of the architecture is provided in Appendix Figure 5.
@@ 4.1 Design
We contribute Jax-RPS, a Jax simulator for the Robotarium. Jax-RPS uses a unicycle dynamics model to faithfully simulate the behavior of the Robotarium’s GRITSBots [23]. To control the GRITSBots, Jax-RPS implements single integrator and unicycle controllers to drive robots to desired positions/poses. To simulate the safety constraints often enforced in real-world deployment, Jax-RPS implements control barrier certificates for collision avoidance, which solve the quadratic program described in [24]. Finally, all functionality in Jax-RPS is designed to be compatible with Jax’s jit compilation for accelerated execution and vmap for parallelization. Overall, this results in Jax-RPS being optimized for data-driven paradigms, such as Multi-Agent RL, while maintaining its faithfulness to real-world conditions and deployment considerations.
@@ 4.2 Evaluation
| Position | RPS | 1425.28 4.79 | N/A |
| Jax-RPS | 9.18 0.10 | 0.00 0.00 | |
| Pose | RPS | 1713.87 5.39 | N/A |
| Jax-RPS | 11.42 0.07 | 0.00 0.00 |
@@ 5 Benchmark Scenarios
Material Transport: Robots with varying speeds and capacities must collaborate to unload two material depots, one further and with less material, and one closer with more material. Task performance is measured by the amount of material remaining to be unloaded.
@@ 6.1 Design
HetMananger provides a user-friendly interface for injecting heterogeneity into multi-robot environments. It supports common heterogeneous representations: robot IDs, class IDs, and capability vectors, and can be extended to include more. It manages the robot’s awareness of heterogeneity, where the observation space can be configured to include any representation.
@@
Scenario interface is a standard structure for defining new tasks within JaxRobotarium, aligned with the structures used in other learning frameworks such as VMAS [14], JaxMARL [11], and MARBLER [12]. At its core, it provides the interface for defining a task-specific observation, reward, and transition function.
@@
We choose to integrate with JaxMARL as MARL framework, given its SOTA performance and extensive baseline algorithm implementations. However, note that any Jax multi-agent learning framework designed for the output structure of the Scenario interface is compatible with JaxRobotarium.
@@ 7.1 Experiment setup
We demonstrate the benchmarking and sim-to-real evaluation capabilities of JaxRobotarium by training 5 policies in parallel for all tasks using the following SOTA MARL approaches:
@@
QMIX [29]: an established multi-agent Q-learning algorithm where decentralized Q-networks predict the Q-values aggregated into a global Q-value by a hypernetwork-generated centralized critic conditioned on the global state.
@@ 7.1.1 Results
Further, we note that independent learning (IPPO) is surprisingly stronger in our multi-robot coordination scenarios compared to its centralized counterpart (MAPPO). We believe this is likely due to the limited extent of partial observability in the benchmark scenarios (e.g., robots can generally observe their neighbors except in Navigation). This suggests that in settings without significant partial observability, independent learning is a simpler and effective alternative to centralized critic approaches, in line with prior results comparing MAPPO and IPPO on StarCraft tasks [30].
@@
| Transport | Success Rate | 1.0 0.0 | 1.0 0.0 | 1.0 0.0 | 1.0 0.0 | 1.0 0.0 | 1.0 0.0 | 0.8 0.18 | 1.0 0.0 |
@@
| (MAPF) | Success Rate | 0.0 0.00 | 1.0 0.0 | 1.0 0.0 | 1.0 0.0 | 0.2 0.18 | 1.0 0.0 | 1.0 0.0 | 1.0 0.0 |
@@
| Foraging | Reward | 6.2 1.07 | 6.2 1.07 | 1.0 0.89 | 6.6 2.41 | 5.2 0.18 | 5.2 0.18 | 2.0 1.07 | 11.0 0.0 |
| Resources Foraged | 1.2 0.18 | 1.2 0.18 | 0.2 0.18 | 1.2 0.45 | 1.0 0.0 | 1.0 0.0 | 0.4 0.22 | 2.0 0.0 | |
@@
Sim2Real: As one would expect, we found reasonable gaps between the simulated and real world runs for most tasks across methods (see Table 4). Some of this gap could be attributed to only having limited real-world runs and the stochasticity of the real world compared to the deterministic simulator. However, two of our new tasks (Predator Prey and Continuous-RWARE) resulted in surprisingly large sim2real performance gaps, particularly for QMIX and PQN. Our qualitative analysis of the rollouts revealed that they tasks contain inflection points that can exacerbate modest differences into large differences in the environment and robot behavior over time. For instance, small positional changes to the robots in Predator-Prey can sometimes elicit qualitatively different prey behaviors which cascade into dramatic differences between the simulated and real-world rollouts. In contrast, existing MARBLER tasks are less sensitive to such variations, as robots mostly interact with a static environment. The observed gaps in the new tasks underscore the importance of hardware evaluation, especially in multi-robot scenarios where task outcomes are closely tied to interaction dynamics. The ability to accurately simulate single robots does not necessarily imply accurate multi-robot simulations.
@@
Domain Randomization: Inspired by the above observations, we applied simple domain randomization through action noise to learn a more robust policy. We found that training with action noise can reduce the sim-to-real gap in Predator Prey for both QMIX and PQN and even improve task performance in simulation (see Table 5).
in at, discovery, foraging, mt, navigation, pp, rware, warehouse

### v3

@@ Abstract
Multi-agent reinforcement learning (MARL) has emerged as a promising solution for learning complex and scalable coordination behaviors in multi-robot systems. However, established MARL platforms (e.g., SMAC and MPE) lack robotics relevance and hardware deployment, leaving multi-robot learning researchers to develop bespoke environments and hardware testbeds dedicated to the development and evaluation of their individual contributions. The Multi-Agent RL Benchmark and Learning Environment for the Robotarium (MARBLER) is an exciting recent step in providing a standardized robotics-relevant platform for MARL, by bridging the Robotarium testbed with existing MARL software infrastructure. However, MARBLER lacks support for parallelization and GPU/TPU execution, making the platform prohibitively slow compared to modern MARL environments and hindering adoption. We contribute JaxRobotarium, a Jax-powered end-to-end simulation, learning, deployment, and benchmarking platform for the Robotarium. JaxRobotarium enables rapid training and deployment of multi-robot reinforcement learning (MRRL) policies with realistic robot dynamics and safety constraints, supporting both parallelization and hardware acceleration. Our generalizable learning interface provides an easy-to-use integration with SOTA MARL libraries (e.g., JaxMARL). In addition, JaxRobotarium includes eight standardized coordination scenarios, including four novel scenarios that bring established MARL benchmark tasks (e.g., RWARE and Level-Based Foraging) to a realistic robotics setting. We demonstrate that JaxRobotarium retains high simulation fidelity while achieving dramatic speedups over baseline (20x in training and 150x in simulation), and provides an open-access sim2real evaluation pipeline through the Robotarium testbed, accelerating and democratizing access to multi-robot learning research and evaluation. Our code is publicly available on github.
@@ 1 Introduction
Despite its potential, learning-based algorithms are seldom deployed and evaluated on physical multi-robot platforms. Even when physical deployments are undertaken, they tend to rely on bespoke and closed platforms and testbeds, limiting reproducibility and benchmarking. In stark contrast, the multi-agent reinforcement learning (MARL) research community outside of robotics benefits from standardized and structured benchmarks, simulation, and learning environments (e.g., Multi-Agent Particle Environment (MPE) [9], StarCraft Multi-Agent Challenge (SMAC) [10], and JaxMARL [11]). However, these environments often abstract away robotics-relevant characteristics (e.g., robot dynamics and collision avoidance) and result in a considerable sim2real gap [8], limiting their ability to faithfully evaluate learning approaches tailored to multi-robot systems.
MARBLER [12] is a recent framework that attempts to meet this critical need. MARBLER acts as a bridge between the standardized infrastructure designed for MARL algorithms and the Robotarium [13] (an open and remotely-accessible multi-robot testbed). However, MARBLER does not support parallelization and execution on GPU/TPU, resulting in prohibitively slow training speeds that hinder adoption and comprehensive evaluations.
@@
JaxRobotarium includes three key contributions. First, we contribute a new Jax-based simulator for the Robotarium that is designed from the ground-up and modeled after the existing Robotarium Python Simulator (RPS). Our new simulator supports parallelization and GPU/TPU execution, enabling dramatic improvements in efficiency by reducing the time and computational overhead typically associated with learning multi-robot coordination in realistic settings. Second, we contribute a new intuitive interface to facilitate integration with existing MARL libraries, such as JaxMARL [11], allowing researchers to leverage and build upon existing algorithmic implementations with minimal setup. Third, we contribute a new unified benchmark for MRRL that currently supports 8 representative multi-robot coordination scenarios, with flexible tools to rapidly create new ones or customize the existing ones. The unique benefits of JaxRobotarium include:
@@
We validate the utility and benefits of JaxRobotarium through extensive efficiency evaluations, algorithm benchmarking, and comprehensive sim2real transfer experiments. We demonstrate that, compared to MARBLER [12], JaxRobotarium simulates robot trajectories up to 150 times faster and trains multi-robot policies up to 20 times faster. We benchmarked four MARL algorithms on the 8 scenarios provided by JaxRobotarium, revealing new insights into the algorithms’ utility across different multi-robot coordination scenarios. We systematically studied the sim2real gap of JaxRobotarium using data from over 200 real-world deployments of multi-robot policies that were entirely trained in simulation. We note that JaxRobotarium is not a panacea for the sim2real gap. Rather, JaxRobotarium helps study and reduce the sim2real gap by enabling rapid training and deployment. Our analysis reveals that while many task-algorithm combinations result in negligible performance differences when deployed on hardware, certain combinations of MARL algorithms and task characteristics exhibit noticeable performance differences. However, we show that such differences can be largely mitigated through simple randomization techniques such as noisy actions. We believe JaxRobotarium will serve as a powerful tool for the multi-robot systems community to advance the development, benchmarking, and sim2real transfer of learning algorithms.
@@ 2 Related Work
The platforms to facilitate learning multi-robot coordination strategies should have the following characteristics: (i) support efficient training and evaluation of policies [11, 14]; (ii) have realistic simulations of robots [8, 13], taking into account their dynamic constraints and the need for lower-level collision avoidance; (iii) have a convenient interface with common MARL algorithms and benchmarking tasks [11, 12, 15]. Centering around these key requirements, we evaluate and compare available MRRL and MARL frameworks in Table 1. The Multi-Agent Particle Environment (MPE) [9] is a widely used framework for evaluating MARL algorithms on a diverse suite of 2D tasks. While MPE was enhanced by successors including VMAS [14], BenchMARL [15], and JaxMARL [11], none of them realistically reflects the dynamics of real robots, impeding straightforward sim2real evaluation of MRRL policies. IsaacGym [16] and IsaacLab [17] offer high-fidelity simulation but are geared toward single-robot learning across varied embodiments. Multi-robot learning in these frameworks requires additional infrastructure for integration with MARL libraries, collision avoidance, benchmark tasks, and a sim2real pipeline, none of which are natively provided. Thus, to better evaluate MRRL policies, researchers must build task-specific test platforms [18, 19, 20] or reproduce previously proposed test beds, such as Cambridge RoboMaster [21]. As a contrastive line of efforts, the Robotarium [13] is a publicly available, multi-robot testbed that supports testing the coordination of up to 20 GRITSBots [22] robots. The Robotarium provides a simulator with implemented collision avoidance using control barrier functions [23] and offers sim2real evaluation through its hardware test bed. It is the only publicly available and free-to-use multi-robot hardware testbed in the world, and has been widely used for multi-robot research in areas including task allocation [24], heterogeneous coordination [25], motion planning [26], and safety [27]. To bridge the gap between MARL and a real multi-robot testbed, [12] contributes MARBLER, which is an open platform that promotes the comprehensive evaluation of MRRL algorithms on the Robotarium. Unfortunately, MARBLER only runs on the CPU without parallelization, resulting in prolonged training and tedious development cycles. JaxRobotarium builds upon the philosophy of MARBLER, marrying it with the computational efficiency offered by Jax and GPU parallelization.
@@
| MARL Platforms (MPE [9], SMAC [10], VMAS [14], | ✓ | ✗ | ✗ | ✗ | ✓ | N/A |
| BenchMARL [15], JaxMARL [11], Google Research Football [28]) | ||||||
| IsaacGym [16], IsaacLab [17] | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ |
@@ 3 JaxRobotarium Platform
JaxRobotarium is designed to achieve the training speeds of Multi-Agent RL in Multi-Robot RL while offering standardized, public, and free sim2real evaluation through the Robotarium. To this end, JaxRobotarium consists of the following key components,
@@
A Jax-based simulator for the Robotarium, compatible with parallelization and GPU/TPU execution.
@@
A direct deployment pipeline for sim2real evaluation on the Robotarium.
We provide a full overview of the architecture in the Appendix (Figure 5).
@@ 4.1 Design
We contribute Jax-RPS, a Jax simulator for the Robotarium. Jax-RPS uses a unicycle dynamics model to faithfully simulate the behavior of the Robotarium’s GRITSBots [29]. To control the GRITSBots, Jax-RPS implements single integrator and unicycle controllers to drive robots to desired positions/poses. To enforce the safety constraints often required in real-world deployment, Jax-RPS implements control barrier certificates for collision avoidance, which solve the quadratic program described in [30]. Finally, all functionality in Jax-RPS is designed to be compatible with Jax’s jit compilation for accelerated execution and vmap for parallelization. Overall, this results in Jax-RPS being optimized for data-driven paradigms, such as Multi-Agent RL, while maintaining its faithfulness to real-world conditions and deployment considerations.
@@ 4.2 Evaluation
| Position | Default | 1425.28 4.79 | N/A |
| JaxRobotarium | 9.18 0.10 | 0.00 0.00 | |
| Pose | Default | 1713.87 5.39 | N/A |
| JaxRobotarium | 11.42 0.07 | 0.00 0.00 |
@@ 5 Benchmark Scenarios
Material Transport: Robots with varying speeds and capacities must collaborate to unload two material depots, one further with less material, and one closer with more material. Task performance is measured by the amount of material remaining to be unloaded.
@@ 6.1 Design
HetMananger provides an interface for injecting heterogeneity into multi-robot environments. It supports common heterogeneous representations: robot IDs, class IDs, capability vectors, and can be extended to include more. It manages the robot’s awareness of heterogeneity, where the observation space can be configured to include any representation.
@@
We choose JaxMARL as the MARL framework, given its SOTA performance and extensive baseline algorithm implementations. However, note that any Jax multi-agent learning framework designed for the output structure of the Scenario interface is compatible with JaxRobotarium.
@@ 7.1 Experiment setup
We demonstrate the benchmarking and sim2real evaluation capabilities of JaxRobotarium by training 5 policies in parallel for all tasks using the following SOTA MARL approaches:
@@
QMIX [35]: an established multi-agent Q-learning algorithm where decentralized Q-networks predict agent Q-values aggregated into a global Q-value by a hypernetwork-generated centralized critic conditioned on the global state.
@@ 7.1.1 Results
Further, we note that independent learning (IPPO) is surprisingly stronger in our multi-robot coordination scenarios compared to its centralized counterpart (MAPPO). We believe this is likely due to the limited partial observability in the configured benchmark scenarios (e.g., robots can generally observe their neighbors except in Navigation). This suggests that in settings without significant partial observability, independent learning is a simpler and effective alternative to centralized critic approaches, in line with prior results comparing MAPPO and IPPO on StarCraft tasks [36].
@@
| Transport | Success Rate | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 0.8 0.18 | 1.0 0.00 |
@@
| (MAPF) | Success Rate | 0.0 0.00 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 | 0.2 0.18 | 1.0 0.00 | 1.0 0.00 | 1.0 0.00 |
@@
| Foraging | Reward | 6.2 1.07 | 6.2 1.07 | 1.0 0.89 | 6.6 2.41 | 5.2 0.18 | 5.2 0.18 | 2.0 1.07 | 11.0 0.00 |
| Resources Foraged | 1.2 0.18 | 1.2 0.18 | 0.2 0.18 | 1.2 0.45 | 1.0 0.0 | 1.0 0.00 | 0.4 0.22 | 2.0 0.00 | |
@@
Sim2Real: As one would expect, we found reasonable gaps between the simulated and real world runs for most tasks across methods (see Table 4). Some of this gap could be attributed to only having limited real-world runs and the stochasticity of the real world compared to the deterministic simulator. However, two of our new tasks (Predator Prey and Continuous-RWARE) resulted in surprisingly large sim2real performance gaps, particularly for QMIX and PQN. Our qualitative analysis of the rollouts revealed that these tasks contain inflection points that can exacerbate modest differences into large differences in the environment and robot behavior over time. For instance, small positional changes to the robots in Predator-Prey can sometimes elicit qualitatively different prey behaviors which cascade into dramatic differences between the simulated and real-world rollouts. In contrast, existing MARBLER tasks are less sensitive to such variations, as robots mostly interact with a static environment. The observed gaps in the new tasks underscore the importance of hardware evaluation, especially in multi-robot scenarios where task outcomes are closely tied to interaction dynamics. The ability to accurately simulate single robots does not necessarily imply accurate multi-robot simulations.
@@
Domain Randomization: Inspired by the above observations, we applied simple domain randomization through action noise to learn a more robust policy. We found that training with action noise can reduce the sim2real gap in Predator Prey for both QMIX and PQN and improve task performance in simulation (see Table 5).